OpenStudy (anonymous):

The simple interest on an investment of $8000 for one year is $360. What is the interest rate? A.0.045% B.0.45% C.4.5% D.45%

OpenStudy (anonymous):

A

OpenStudy (anonymous):

how do you know it is A. can you explain?

OpenStudy (ranga):

Not A.

OpenStudy (anonymous):

ok then what would it be

OpenStudy (texaschic101):

I = PRT (interest = principle * rate * time)...but since we are looking for R. lets change the formula around..... I / PT = R I = 360 P = 8000 T = 1 Now we sub 360/(8000)(1) = R 360/8000 = R 0.045 = R....change to a percent = 4.5% (C)

OpenStudy (anonymous):

oh thank you. i like how you explained it. thank you for your help

OpenStudy (texaschic101):

glad to help :)